What are the values of x in the equation x(x-6)=4(x+6)
Answer by ewatrrr(24785)


Hi
x(x-6)=4(x+6)
x^2 - 6x = 4x + 24
x^2 -10x - 24 = 0
(x-12)(x+2) = 0, x = -2, 12